### v1

# DecompDreamer: Advancing Structured 3D Asset Generation with Multi-Object Decomposition and Gaussian Splatting

Abstract
Text-to-3D generation saw dramatic advances in recent years by leveraging Text-to-Image models. However, most existing techniques struggle with compositional prompts, which describe multiple objects and their spatial relationships. They often fail to capture fine-grained inter-object interactions. We introduce DecompDreamer, a Gaussian splatting-based training routine designed to generate high-quality 3D compositions from such complex prompts. DecompDreamer leverages Vision-Language Models (VLMs) to decompose scenes into structured components and their relationships. We propose a progressive optimization strategy that first prioritizes joint relationship modeling before gradually shifting toward targeted object refinement. Our qualitative and quantitative evaluations against state-of-the-art text-to-3D models demonstrate that DecompDreamer effectively generates intricate 3D compositions with superior object disentanglement, offering enhanced control and flexibility in 3D generation. Project page : https://decompdreamer3d.github.io
Figure 1: Illustration of high-quality, complex 3D scenes generated by DecompDreamer for prompts containing 6-11 objects. Existing methods struggle to generate all objects accurately and fail to maintain correct spatial relationships.
footnotetext: * Equal Contribution.footnotetext: $\dagger$ Corresponding author: unath@asu.edu
1 Introduction
Figure 2: (a) Illustration of feature-object mismatch. The image generated by StableDiffusion 2.1 using the prompt ”A blue bird sitting on a black branch of a white tree” (b) 3D representations generated by LucidDreamer [20] and DecompDreamer.
3D representations play a crucial role in many applications including perception and autonomy, as well as in the media industries, enabling the creation of immersive and realistic digital environments. These industries strive to replicate real-world complexity by incorporating high-fidelity 3D objects that interact dynamically within virtual spaces. Real-world scenes consist of objects and agents with complex spatial relationships and potential interactions, with their surroundings, that can only be described by intricate semantic relationships. Recent methods employ 2D lifting techniques [6, 21, 24, 32, 40] to generate 3D models from text using pre-trained diffusion models [34, 29, 35] and Score Distillation Sampling (SDS) [32]. While effective for single objects with simple text prompts, these methods struggle with compositional prompts involving multiple objects, with quality significantly degrading as inter-object relationships increase. This decline is primarily due to (1) feature-object mismatch and (2) poor modeling of 3D spatial relationships.
Feature-object mismatch arises as a fundamental limitation of text-image embeddings [33]. As text complexity increases, matching attributes to their corresponding objects becomes increasingly challenging. For instance, in the prompt “A blue bird sitting on a black branch of a white tree” (Figure 2(a)), the text-to-image model confuses the attributes “blue” and “white,” resulting in a bird with blue-white coloration instead of the intended output. Furthermore, as the number of inter-object relationships grows, existing methods fail to maintain spatial coherence, leading to incorrect placements, overlapping artifacts, or missing contextual dependencies. For example, in Figure 2(b), the generated 3D representation by LucidDreamer [20] deteriorates as the complexity and number of inter-object relationships increase.
While generating a high-fidelity 3D representation corresponding to compositional text prompts is crucial, it is equally important to break the scene down into smaller, well-defined components. A monolithic 3D output limits flexibility, making it difficult to refine individual objects or modify their attributes without affecting the entire scene. A decomposed representation provides greater control, allowing artists and developers to tweak specific parts while preserving overall scene coherence.
In this paper, our goal is to generate decomposable 3D representations corresponding to complex text prompts involving intricate inter-object relationships. Let us consider how an artist might approach this problem [9]. An expert 3D modeler tasked with designing a 3D scene consisting of multiple objects would likely begin by identifying the number of simple objects, estimating their approximate relative positions, and then creating meshes for each object while refining their spatial relationships. Following this simple yet effective strategy, we propose DecompDreamer, a Gaussian-splatting-based optimization routine that decomposes complex 3D representations into smaller components and progressively refines them. Our approach is built on three core design choices:
Structured 3D Decomposition and Initialization: Given a text prompt, we use a Vision-Language Model (VLM) to decompose the scene into individual objects, infer their spatial arrangement, and establish inter-object relationships. Each object is represented using 3D Gaussians, enabling structured parameter initialization via coarse 3D point clouds generated by Point-E [28]. Unlike implicit representations such as NeRF [25], which lacks object-level control, this approach ensures a spatially consistent and well-structured starting point.
Progressive Optimization: Existing SDS-based methods optimize entire scenes at once [6, 21, 24, 32, 40, 43] or iteratively refine objects, relationships and full scene [13, 14]; often leading to entangled representations and inconsistencies. While effective for simple objects or those with few inter-object relationships, their performance degrades significantly as the number of relationships increases. In this paper, we propose a progressive optimization strategy inspired by artistic workflows. Unlike existing methods, we do not primarily focus on optimizing the entire scene at once. Instead, we emphasize modeling inter-object relationships and refining individual objects, with occasional scene-level optimization. We start with joint optimization of objects with respect to their relationships, then gradually shift focus toward individual object refinement. Optimizing objects in a compositional environment is significantly more challenging than in isolation as it requires accounting for spatial dependencies and relative viewing direction. To address this, we incorporate object-view-aware optimization, improving object disentanglement and ensuring a coherent 3D structure. Once the scene is well-formed, we apply targeted optimization within the relational context to refine fine-grained object details while preserving spatial coherence.
Flow Matching Distillation: To effectively model objects and their relationships, we extend diffusion-based SDS to flow-based SDS, leveraging rectified flow models [1, 2] for smoother and more controlled optimization. Flow matching enables stable and efficient optimization by learning a direct vector field, eliminating stochastic noise and improving convergence.
With these advancements, DecompDreamer is capable of generating high-quality compositional 3D objects (Figure 1, 4) while achieving enhanced disentanglement for individual objects (Figure 5). We evaluate DecompDreamer through both qualitative and quantitative analyses. A user study indicates that DecompDreamer outperforms existing models, including Gala3D [43], GraphDreamer [13], LucidDreamer [20], MVDream [36], and HiFA [45], in terms of image quality and semantic alignment. Additionally, we perform quantitative evaluations using CLIP score [33], and Pick-A-Pic [19] on 30 compositional text prompts containing 2 to 11 objects, demonstrating DecompDreamer’s superior performance in compositional 3D generation.
Our contributions can be summarized as:
  1. 1.
    We introduce DecompDreamer, a 3D Gaussian-based training routine that decomposes complex 3D representations into smaller, manageable parts and applies progressive optimization to refine relationships and individual objects, using a novel flow matching distillation approach.
  2. 2.
    DecompDreamer achieves strong disentanglement across individual object components using view-aware optimization, enhancing interpretability and controllability.
  3. 3.
    We extensively validate our approach using a diverse set of complex text prompts and benchmark it both qualitatively and quantitatively against several state-of-the-art 3D generators [21, 6, 40, 13, 43]. Empirical evaluations demonstrate that DecompDreamer consistently produces well-structured, high-fidelity compositional 3D objects.
2 Related Work
Text-to-3D Generation converts textual descriptions into 3D objects and is broadly categorized into 3D supervised and 2D lifting techniques. 3D supervised methods  [8, 17, 28, 41] use paired text and 3D data for training, enabling high-quality 3D generation. However, they struggle to generalize to compositional scenes, as most 3D datasets  [10, 23] focus on single-object models. In contrast, 2D lifting techniques  [32, 21, 6, 40, 44, 16, 7] refine 3D representations using pre-trained 2D diffusion models. DreamFusion  [32] introduced Score Distillation Sampling (SDS) to align 3D content with a 2D diffusion prior, inspiring further advancements  [20, 21, 39, 38, 44]. However, these methods struggle with compositional 3D generation as they do not model inter-object relationships individually but instead treat the entire scene as a single entity, leading to poor disentanglement of objects and relationships.
To address this problem, recent methods [11, 4, 31, 42] incorporate 3D layouts with non-overlapping bounding boxes. While effective for structured scenes, they struggle with overlapping objects. GraphDreamer [13] addresses this issue by using a scene graph to represent 3D scenes, separately modeling objects and their relationships. However, it relies on position embeddings to map 3D locations to objects, making optimization challenging. Consequently, it struggles to scale to scenes with a large number of objects and suffers from poor disentanglement. Gala3D [43] enables compositional 3D generation by optimizing Gaussians for individual objects and the overall scene. However, it struggles with complex interactions, particularly overlapping objects, leading to spatial inaccuracies. CompGS [14] enhances initialization using pre-trained text-to-3D models yet struggles with intricate relationships. We propose DecompDreamer to address these challenges and enable the generation of compositional scenes with a large number of objects and intricate relationships. Similar to Gala3D and CompGS, we use Gaussians as the 3D representation. However, instead of jointly optimizing objects and the entire scene, we adopt a progressive optimization approach, transitioning from inter-object relationships to individual object refinement for more structured and disentangled 3D generation.
Rectified Flow Models (RFM)  [3, 22] have recently gained attention as an alternative generative approach to diffusion models. Unlike traditional diffusion models  [15, 37], which rely on stochastic sampling via a stepwise denoising process, RFMs directly learn smooth, continuous transport between data distributions, eliminating iterative noise perturbations and leading to more efficient sampling and improved generation quality. Recent works  [12] have demonstrated the effectiveness of RFMs for large-scale image and video generation. In this paper, we apply RFMs to 3D generation tasks, showcasing their capabilities.
3 Preliminaries
Gaussian Splatting  [18] models a scene with a collection of 3D Gaussians, characterized by means, covariances, colors, and opacities. These Gaussians are projected onto the image plane as 2D Gaussians for rendering, where colors are composited using alpha blending in a front-to-back order. Compared to NeRF-based methods  [5, 26], Gaussian Splatting offers higher reconstruction fidelity, faster rendering, and lower memory consumption during training. The rendering process can be formulated as $r=g(\theta,c)$, where $\theta$ denotes the 3D Gaussians, $g$ represents the renderer, $c$ captures the camera parameters, and $r$ is the resulting image.
Score Distillation Sampling (SDS)  [32] optimizes a 3D representation $\theta$ using a pre-trained 2D diffusion model $\phi$. The objective is to refine the 3D representation so that images rendered from various camera projections closely match those generated by the pre-trained 2D text-to-image diffusion model. The optimization process is guided by the SDS gradient, formulated as:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(z;y)=\mathbb{E}_{t,\epsilon}\Big{[}\omega(t)\big{(}\hat{\epsilon}_{\phi}(z_{t},t,y)-\epsilon\big{)}\frac{\partial z}{\partial\theta}\Big{]},$ (1)
where, $\hat{\epsilon}_{\phi}$ represents the noise predicted by the model $\phi$, and $z_{t}$, $t$, and $y$ correspond to the noisy image, timestep, and text embedding, respectively. $\epsilon$ denotes Gaussian noise, and $\omega(t)$ is the weighting function applied at each timestep. SDS treats $\hat{\epsilon}_{\phi}(.)$ as a score function that estimates the sampled noise $\epsilon$ present in $z_{t}$ at a specific noise level $t$, conditioned on the text embedding $y$. The optimization minimizes the discrepancy between the predicted noise and the actual noise, guiding the 3D representation $\theta$ to produce renderings that align with the given text prompt.
Figure 3: Overall pipeline of DecompDreamer. Given a text prompt, a VLM generates an image and a scene graph. The scene graph identifies objects and their relationships, while the image provides spatial structure. Using this information, we perform coarse initialization followed by a progressive optimization strategy. Initially, optimization focuses on joint relationship modeling before gradually shifting towards targeted and view-aware object optimization, resulting in high-fidelity 3D objects with strong object disentanglement.
4 Method
Our objective is to generate high-quality 3D assets for complex text prompts that describe intricate inter-object relationships. Figure 3 shows an overview, with details described below. In section  4.1, we discuss rectified flow based distillation. In section  4.2, we discuss the structured decomposition of compositional text into objects and their relationship via VLM. Finally, in section  4.3, we discuss the progressive optimization process, transitioning from relationship modeling to object refinement for high-fidelity compositional 3D generation.
4.1 Flow Matching Distillation (FMD)
Rectified flow models employ a linear interpolation forward process given by $z_{t}=(1-\sigma_{t})z+\sigma_{t}\epsilon$, or equivalently, $z_{t}=z-\sigma_{t}(\epsilon-z)$. This process interpolates between the data $z$ and the noise $\epsilon$ using a time-dependent weighting function $\sigma_{t}$, which is sampled from a uniform distribution $\mathcal{U}(0,1)$. For simplicity, we denote $(\epsilon-z)$ by $u(z,\epsilon)$, representing the actual vector field.
FMD optimizes a 3D representation $\theta$ using a pre-trained rectified flow-based model $\phi$. The optimization, which closely aligns with the SDS gradient, is formulated as
$\nabla_{\theta}\mathcal{L}_{\text{FMD}}(z;y)=\mathbb{E}_{t,\epsilon}\Big{[}\sigma_{t}\big{(}v_{\phi}(z_{t},t,y)-u(z,\epsilon)\big{)}\frac{\partial z}{\partial\theta}\Big{]},$ (2)
where $v_{\phi}(.)$ is the vector field predicted by the model $\phi$, and where $z_{t}$, $t$, and $y$ correspond to the noisy image, the timestep, and the prompt embedding, respectively. This optimization minimizes the discrepancy between the predicted and the actual vector fields, thereby guiding the 3D representation $\theta$ to produce renderings that are consistent with the given text prompt.
4.2 Structured Decomposition with VLMs
Before designing a complex 3D asset, an artist typically begins with careful planning and wireframing the overall structure. This involves determining the placement, size, and orientation of individual objects within the scene. Inspired by this process, we leverage a VLM, specifically ChatGPT-4o [30], to structure our 3D representation. Given a text prompt $t^{g}$, the VLM generates an image $I$ along with a corresponding scene graph $G=(V,E)$, where $V$ represents the set of objects and $E$ captures the inter-object relationships. We explicitly instruct the VLM to generate a connected graph, ensuring each node has at least one edge.
The objects identified by the VLM from the text prompt $t^{g}$ are denoted as $\{O_{i}\}_{i=1}^{n}$, where $n$ is the total number of objects extracted. These objects serve as the nodes $V=\{v_{i}\}_{i=1}^{n}$ in the scene graph $G$. The VLM extracts additional descriptions in the form of attributes $\{a_{ij}\}_{j=1}^{m}$ from the text prompt $t^{g}$ and associates them with their corresponding nodes in the scene graph. By concatenating an object $O_{i}$ with its attributes, we construct an object prompt $t^{o}_{i}$. For example, in Figure 3, the combination of the object and its attributes yields the object prompt: “A milkshake with two straws”. Similarly, the relationships between object pairs $(O_{i},O_{j})$, as described in $t^{g}$, are mapped to edges $e_{ij}$ within the scene graph $G$. These relationships capture the contextual interaction between the objects. To represent these interactions, we construct edge prompts $t^{e}_{ij}$ by concatenating the involved objects and the relationship attribute. By integrating both object-level and relational details through object and edge prompts, the scene graph serves as a comprehensive blueprint, effectively guiding the 3D representation process.
Additionally, we pass the generated image $I$ and scene graph $G$ to the VLM to infer spatial and geometric properties such as center coordinates, size, and orientation of each object. These properties collectively define the spatial arrangement of the 3D representation.
Coarse Initialization. Our primary objective is to accurately capture complex object relationships while ensuring proper disentanglement of individual elements within the scene. To achieve this, we initialize the scene using 3D Gaussians, where each Gaussian is assigned to a specific object $O_{i}$. The Gaussians are initialized from point clouds generated by Point-E [28] based on object-specific prompts. Each object’s point cloud is aligned using its inferred center, scale, and orientation to maintain scene consistency. The full scene representation is obtained by aggregating all aligned point clouds, forming a structured initialization for further optimization. For efficient disentangled optimization, we maintain a mapping that tracks Gaussians corresponding to each object, enabling independent refinement of properties such as mean, covariance, and opacity. Further details, including transformation equations and formal definitions, can be found in Section A.3 of the Appendix.
4.3 Progressive FMD Optimization
We begin by optimizing objects with respect to inter-object relationships before gradually shifting to targeted object refinement. Section 4.3.1 details two strategies for modeling inter-object interactions, while Section 4.3.2 addresses view-aware object modeling for disentanglement. Finally, Section 4.3.3 outlines the full optimization pipeline, integrating relationship modeling and object refinement for a coherent 3D representation.
4.3.1 Ensuring Coherent Inter-Object Interactions
In our framework, inter-object relationships are represented using edge prompts, which describe the contextual interactions between object pairs. To effectively model these relationships, we employ two complementary strategies for optimizing the associated Gaussians:
Joint Optimization of Related Objects: Given an edge prompt $t^{e}_{ij}$ describing the relationship between objects $O_{i}$ and $O_{j}$, we jointly optimize the Gaussians corresponding to both objects. This ensures that the spatial and semantic relationships between the objects are preserved in the 3D scene. The FMD loss for this joint optimization is :
$\mathcal{L}_{joint}(O_{i},O_{j},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}},\mathcal{G}_{o_{j}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij}),$ (3)
where, $\mathcal{L}_{\text{FMD}}$ is defined in (2), $\mathcal{G}_{o_{i}}$ and $\mathcal{G}_{o_{j}}$ represent the Gaussians for objects $O_{i}$ and $O_{j}$, respectively, and $y^{e}_{ij}$ is the text embedding corresponding to the edge prompt $t^{e}_{ij}$.
Targeted Optimization for a Single Object: While joint optimization ensures cohesive inter-object relationships, it can sometimes result in undesirable blending, where parts of one object flow into another. To mitigate this, we employ targeted optimization based on the edge prompt, focusing the optimization solely on one object while keeping the other fixed. The goal is to adjust the selected object such that it adheres to the overall relationship without affecting the structural integrity of the other object. The FMD loss for targeted optimization is formulated as:
$\mathcal{L}_{target}(O_{i},y^{e}_{ij})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij}),$ (4)
where, only the Gaussians of object $O_{i}$ are optimized based on the edge prompt $t^{e_{ij}}$.
Scene-Aware Gaussian Translation: When modeling inter-object relationships, it is important to account for errors in object center predictions by the VLM. To address this, we introduce object-specific translation parameters $T^{o}_{i}=(x_{i},y_{i},z_{i})$ for each object $O_{i}$. Before optimizing edges using (3) or (4), Gaussians are translated using these parameters to correct misalignments and ensure accurate spatial relationships. The transformation is applied as:
$\mathcal{G}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i},\quad\mathcal{G}_{o_{j}}\leftarrow\mathcal{G}_{o_{j}}+T^{o}_{j}$
where $T^{o}_{i}=(x_{i},y_{i},z_{i})$ and $T^{o}_{j}=(x_{j},y_{j},z_{j})$ represent the translation adjustments for objects $O_{i}$ and $O_{j}$, respectively.
4.3.2 View-Aware Disentangled Object Optimization
DecompDreamer promotes object disentanglement by assigning each Gaussian to a single object, ensuring clear boundaries. However, solely optimizing Gaussians based on inter-object relationships can lead to object blending and suboptimal refinement of individual objects. This section discusses optimizing objects based on attributes and enhancing disentanglement using negative prompts.
Optimizing Objects within Compositional Environment: A naïve approach would apply SDS-based optimization directly using object prompts, but in a multi-object scene, relative object orientation significantly impacts perception. DreamFusion [32] demonstrated that view-dependent text prompts improve SDS-based optimization by aligning prompts with the camera’s viewpoint. Let $t$ be the base text prompt. The view-dependent embedding $y_{a}$ is generated using $E(t,a)$, which integrates $t$ with the azimuth angle $a$ to produce a view-aware embedding. This method interpolates between front, side, and back views based on $a$, aligning text prompts with the camera’s viewpoint. This alignment enhances orientation-specific details, leading to more coherent 3D outputs and has become standard in optimization-based 3D generation. However, global view-dependent prompts fail in multi-object scenes where a single camera view may not accurately represent all objects. To address this, we introduce object-view-dependent text prompts, ensuring each object’s orientation aligns with its correct perspective. First, each object $O_{i}$ is rendered at a canonical front-facing orientation ($0^{\circ}$ azimuth) to establish a reference view. The VLM then estimates the angular difference $\phi_{i}$ between the current view and the object’s front view, correcting the azimuth as $a_{i}=a-\phi_{i}$ (a sample VLM workflow is provided in Section A.5 of the appendix). The adjusted azimuth $a_{i}$ is incorporated into the object’s text prompt, leading to the object-specific FMD loss:
$\mathcal{L}_{obj}(O_{i},y^{o}_{a_{i}})=\nabla_{\mathcal{G}_{o_{i}}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{i}}),$ (5)
where $y^{o}_{a_{i}}=E(t^{o}_{i},a)$. This adjustment ensures more coherent 3D reconstructions by aligning object prompts with their true viewpoints.
Negative Prompt for Disentangled Object: As discussed earlier, optimizing inter-object relationships can lead to object blending, where one object undesirably flows into another. To mitigate this, we incorporate negative prompts during individual object optimization. The negative prompt for an object $O_{i}$ is constructed by aggregating all objects directly connected to $O_{i}$ in the scene graph, ensuring that the model explicitly avoids incorporating features from these related objects. This approach helps maintain clear object boundaries and preserves structural integrity within the scene.
Figure 4: Qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20] and Magic3D [21]. More results can be found in Section A.6 of the appendix.
4.3.3 Overall Optimization
After initialization, we begin by modeling inter-object relationships before gradually shifting focus toward individual objects. The optimization process is divided into two stages:
1. Stage One: We emphasize joint modeling of inter-object relationships. For each edge in the scene graph, we apply joint optimization (3), while progressively increasing the weight of individual object optimization (5).
2. Stage Two: The focus shifts toward refining individual objects while maintaining inter-object consistency. Here, we perform targeted optimization for each object (4) along with individual object optimization (5). This transition allows fine-grained details to be added while preserving relationships established earlier. The effective FMD loss is:
$\mathcal{L}_{\text{loss}}=\begin{cases}\begin{array}[]{l}\mathcal{L}_{joint}(O_{i},O_{j},y^{e}_{ij})\\ +\lambda.\left(\frac{t}{T}\right)^{2}\sum\limits_{O\in\{O_{i},O_{j}\}}\mathcal{L}_{obj}(O,y^{o}_{a_{o}}),\\ \end{array}&t<\gamma.T\\[8.0pt] \\ \begin{array}[]{l}\sum\limits_{O\in\{O_{i},O_{j}\}}(\mathcal{L}_{target}(O,y^{e}_{ij})\\ +\mathcal{L}_{obj}(O,y^{o}_{a_{o}})),\end{array}&t>\gamma.T\end{cases}$ (6)
where $(i,j)\in E$, $t$ represents the current iteration, $T$ is the total number of iterations, and $\gamma$ controls the transition point between joint and targeted optimization. After a complete cycle of optimizing each edge, we refine the entire scene based on the original text prompt. The full pipeline, including algorithmic details, is provided in Section A.1 of the Appendix.
5 Experiments
Figure 5: (a) Qualitative comparison of object disentanglement in GALA3D, GraphDreamer, and DecompDreamer. Each method displays the full 3D object (left) and its individual components (right). (b) Ablation study on with and without the azimuth adjustment for object-view aware object optimization.
Figure 6: Results of an ablation study on DecompDreamer with the text prompt: “A knight in full shining armor galloping on a brown horse, wearing a blue hat, and holding a sword”. Additional ablation studies can be found in Figure  8 of the appendix.
Figure 7: Qualitative comparison of GALA3D and GraphDreamer against DecompDreamer with SD2.1 and SD3.5 as backbones.
We implement our guidance model using the publicly available diffusion model, stabilityai/stable-diffusion-3.5-medium, with a guidance scale of 7.5. For relationship optimization, we adopt the view-dependent prompting technique from DreamFusion, while for object optimization, we introduce object-view-dependent text prompts, where the corrected azimuth angle is estimated by the VLM. Our 3D Gaussian Splatting rendering pipeline is implemented in PyTorch, with the total number of iterations determined by the number of objects, allocating 1500 iterations per object. Each object’s Gaussian representation is initialized with 80K points. Generating a 3D compositional asset with two objects typically requires 90 minutes on an A100 GPU, averaging 45 minutes per object. Further implementation details and hyperparameter settings can be found in Section A.2 of the Appendix . We qualitatively and quantitatively compare DecompDreamer with state-of-the-art text-to-3D generators on compositional text prompts in Section 5.1. We present our ablation studies in Section 5.2 on significance of various component of DecompDreamer.
5.1 Comparison to State-of-the-art Methods
We qualitatively and quantitatively evaluate DecompDreamer against state-of-the-art (SOTA) text-to-3D models [43, 13, 36, 20, 21] using complex compositional prompts. Another notable work, CompGS [14], also achieves SOTA results, however, we are unable to compare with them as their code is not available at the time of writing this paper. As shown in Figures 1, 4 and 15, DecompDreamer effectively models intricate inter-object relationships while scaling to a large number of objects. GALA3D is capable of generating large-scale environments (e.g., a bedroom) but struggles with modeling relationships, even in simpler cases like a knight riding a horse. This limitation arises because it does not explicitly model individual relationships, relying instead on global scene optimization. GraphDreamer jointly optimizes objects and relationships but suffers from scalability issues, as mapping each 3D point to its corresponding object becomes increasingly difficult. In our experiments, it failed to generate reasonable results for prompts with more than four objects; to evaluate larger scenes, we combined multiple objects. LucidDreamer, MVDream, HiFA, and Magic3D perform well on simple prompts but struggle as object count and complexity increase. Their quality significantly declines in highly compositional scenes. In addition to qualitative analysis, we quantitatively compare SOTA methods using CLIP score [33], Pick-A-Pic [19], and a user study on 30 compositional prompts (2–11 objects) in Table 1. In the user study, 20 participants selected the best matches for each prompt based on semantic alignment and quality, with 54% favoring DecompDreamer. DecompDreamer achieves the highest scores across all three metrics, demonstrating its effectiveness in compositional 3D generation.
Disentanglement in Multi-Object 3D Generation Beyond generating compositional 3D scenes, producing high-quality, disentangled individual objects is equally crucial. Figure 5 compares individual objects from DecompDreamer, GALA3D, and GraphDreamer on complex prompts. Both GALA3D and GraphDreamer exhibit poor disentanglement, with GraphDreamer being more prone to object blending due to its stronger focus on relationship modeling. DecompDreamer achieves superior disentanglement through targeted relationship optimization and view-aware object refinement. Additional results are provided in Section A.4 of the appendix.
Table 1: Comparison of CLIP Score, Pick-A-Pic, and User Study results across different text-to-3D methods. All values are reported as percentages.
Method CLIP Score $\uparrow$ Pick-A-Pic $\uparrow$ User Study $\uparrow$
Magic3D 30.77 9.1 0
HiFA 29.55 16.9 2.4
GraphDreamer 28.07 11.6 18.3
GALA3D 31.07 8.6 10.6
LucidDreamer 32.34 15.7 3.3
MVDream 31.71 15.5 11.4
DecompDreamer (Ours) 34.5 22.5 54
5.2 Ablation Study
Training Routine. We conduct an ablation study to assess the impact of different components in our training routine. Figure 6 compares several variations: ”Base” directly trains on the input prompt without decomposition using SD3.5, failing to generate a coherent 3D structure. $\mathcal{L}_{joint}$ optimizes only relationships, forming a well-structured scene but with poor disentanglement, causing objects to blend. $\mathcal{L}_{joint}+\sum\mathcal{L}_{obj}$ assigns equal weight to relationships and objects, however it confuses the model, leading to structural inconsistencies. For instance, a knight riding a horse appears structurally different from a standalone knight, confusing the model and resulting in poor relationship modeling. In contrast, our progressive optimization strategy, which prioritizes relationships before refining individual objects, significantly enhances structural integrity. The first stage ensures well-formed relationships, while the second stage improves object quality while maintaining these relationships. Additional results and analysis are in Section A.4 of the appendix.
Azimuth Adjustment. We conduct an ablation study to evaluate the impact of azimuth adjustment and object-view-dependent text prompts in (5). Without azimuth correction, objects exhibit noticeable disorientation. As shown in Figure 9, omitting object-view adjustments in the robot example results in misplaced features, such as eyes appearing on the back of its head. This misalignment occurs because the global front view differs from the object’s actual orientation, causing $\mathcal{L}_{obj}$ and $\mathcal{L}_{joint}$ to optimize in conflicting directions. Further results and analysis are in Section A.4 of the appendix.
Comparison with SD2.1-based DecompDreamer. To evaluate DecompDreamer’s optimization routine, we replace the FMD-based Stable Diffusion 3.5 with SDS-based Stable Diffusion 2.1. Compared to other SOTA methods, DecompDreamer with SD2.1 achieves better inter-object relationship modeling and successfully generates all objects in a scene. As shown in Figure 13, under the “hero dragon” prompt, methods like GraphDreamer and MVDream fail to generate all objects, while GALA3D struggles with relationship modeling (e.g., the dragon faces the wrong direction). In contrast, DecompDreamer with SD2.1 correctly captures both objects and their relationships. Further improvements are achieved by replacing SDS with Flow Matching Distillation (FMD), enhancing sharpness and reducing iteration requirements by 75%. Additional discussion are in Section A.4 of the appendix.
6 Conclusion
In this paper, we introduce DecompDreamer, a Gaussian splatting-based text-to-3D generation framework designed to handle compositional prompts with a large number of objects and intricate inter-object relationships. Unlike existing methods, DecompDreamer not only generates high-quality compositional 3D structures but also ensures coherent individual object generation, providing greater flexibility and control for downstream applications. Extensive experiments on prompts containing 2–11 objects validate DecompDreamer’s effectiveness in capturing complex relationships while maintaining object-level disentanglement. While our primary focus has been on modeling inter-object relationships, the generation quality of individual objects is still constrained by SDS optimization, making it susceptible to artifacts such as the Janus problem. However, recent advancements such as MT3D [27] and DreamControl [16] can be seamlessly integrated into our approach to enhance object fidelity if needed. Additionally, DecompDreamer is tailored for inter-object relationships and may not generalize well to decomposing a single entity into finer-grained components. Addressing these challenges may require a more structured, hierarchical representation for single-object decomposition, which we leave for future exploration.
Appendix A Appendix
A.1 Notations and Algorithm
Table 2: Summary of Notations Used in the Algorithm
Notation Description
$t^{g}$ Compositional text prompt
VLM Vision-Language Model
$T_{\text{max}}$ Total number of optimization iterations
$\mathcal{I}$ Generated image representation
$G$ Scene graph extracted from VLM
$N$ Number of objects in the scene
$E$ Edge list (connections between objects)
$O$ Set of all objects in the scene
$t^{o}_{i}$ Text prompt for object $O_{i}$
$t^{e}_{ij}$ Text prompt for edge $(i,j)$
$(x_{i},y_{i},z_{i})$ 3D position of object $O_{i}$
$s_{i}$ Scale factor for object $O_{i}$
$\theta_{i}$ Rotation parameters for object $O_{i}$
$\mathbf{R}(\theta_{i})$ Rotation matrix for object $O_{i}$
$\mathbf{T}(x_{i},y_{i},z_{i})$ Translation matrix for object $O_{i}$
$P(O_{i})$ Point cloud representation of object $O_{i}$
$P(O_{i})_{\text{aligned}}$ Aligned point cloud of $O_{i}$ after transformations
$\mathcal{G}_{o_{i}}$ Gaussian representation of object $O_{i}$
$\mathcal{G}$ Combined Gaussian representation of the scene
$n_{e}$ Number of edges in the scene graph ($|E|$)
$\mathcal{L}_{\text{joint}}$ Loss for Joint Optimization
$\mathcal{L}_{\text{obj}}$ Object-specific FMD loss
$\mathcal{L}_{\text{scene}}$ Scene-level loss
$\mathcal{L}_{\text{target}}$ Loss for targeted Optimization
$\mathcal{L}_{\text{FMD}}$ Flow Matching Distillation loss
$\mathcal{L}_{\text{loss}}$ Total loss used for optimization
$\lambda$ Weight factor to control contribution of $\mathcal{L}_{obj}$
$\gamma$ Weight controlling the phase transition in overall optimization
$z_{ij}$ Rasterized and encoded representation of gaussians
Adam() Adam optimizer for updating Gaussian parameters
DensifyAndPrune() Function to densify and prune Gaussian representations
IsRefinementIteration($t$) Function to check if densification is needed at iteration $t$
Algorithm 1 Structured Decomposition and Progressive FMD Optimization
Input: $t^{g}$: Compositional text prompt, VLM: Vision-Language Model, $T_{\text{max}}$: Total iterations
1:Stage 1: Structured Decomposition and Initialization
2:$\mathcal{I},G,N\leftarrow$ VLM($t^{g}$) $\triangleright$ Generate image and scene graph
3:$E,O,\{t^{o}_{i},t^{e}_{ij},x_{i},y_{i},z_{i},s_{i},\theta_{i}\}\leftarrow$ VLM($\mathcal{I},G$)
4:for $i=1$ to $N$ do $\triangleright$ Convert objects to Gaussians
5:     $P(O_{i})\leftarrow$ Point-E($t^{o}_{i}$)
6:     $P(O_{i})_{\text{aligned}}\leftarrow s_{i}\cdot\mathbf{R}(\theta_{i})\cdot P(O_{i})+\mathbf{T}(x_{i},y_{i},z_{i})$
7:     $\mathcal{G}_{o_{i}}\leftarrow P(O_{i})_{\text{aligned}}$
8:end for
9:$\mathcal{G}=\bigcup_{i=1}^{N}\mathcal{G}_{o_{i}}$triangleright Unified Gaussian representation
10:Stage 2: Progressive FMD Optimization
11:$n_{e}\leftarrow|E|$triangleright Number of edges
12:for $t=1$ to $T_{\text{max}}$ do
13:     if $t<\gamma.T_{\text{max}}$ then $\triangleright$ Joint Optimization
14:         if $t\bmod(n_{e}+1)\neq 0$ then
15:              $i,j\leftarrow E[t\bmod n_{e}]$
16:              $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i},\quad\mathcal{G}^{\prime}_{o_{j}}\leftarrow\mathcal{G}_{o_{j}}+T^{o}_{j}$
17:              $\mathcal{L}_{\text{joint}}\leftarrow\nabla_{\mathcal{G}^{\prime}_{o_{i}},\mathcal{G}^{\prime}_{o_{j}}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij})$
18:              $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{O}})$
19:              $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{joint}}+\lambda\left(\frac{t}{T_{\text{max}}}\right)^{2}\mathcal{L}_{\text{obj}}$
20:         else
21:              for $i=1$ to $N$ do
22:                  $\mathcal{G}^{\prime}_{o_{i}}\leftarrow\mathcal{G}_{o_{i}}+T^{o}_{i}$
23:              end for
24:              $\mathcal{L}_{\text{scene}}\leftarrow\nabla_{\mathcal{G}}\mathcal{L}_{\text{FMD}}(z;y^{g})$
25:              $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{scene}}$
26:         end if
27:     $\triangleright$ Targeted Optimization
28:         $i,j\leftarrow E[t\bmod n_{e}]$
29:         $\mathcal{L}_{\text{target}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{e}_{ij})$
30:         $\mathcal{L}_{\text{obj}}\leftarrow\sum\limits_{O\in\{O_{i},O_{j}\}}\nabla_{\mathcal{G}_{O}}\mathcal{L}_{\text{FMD}}(z;y^{o}_{a_{O}})$
31:         $\mathcal{L}_{\text{loss}}\leftarrow\mathcal{L}_{\text{target}}+\mathcal{L}_{\text{obj}}$
32:     end if
33:     $\mathcal{G}\leftarrow$ Adam($\nabla\mathcal{L}_{\text{loss}}$) $\triangleright$ Update step
34:     if IsRefinementIteration($t$then
35:         $\mathcal{G}\leftarrow$ DensifyAndPrune($\mathcal{G}$)
36:     end if
37:end for
A.2 Hyperparameters
Our optimization framework follows the learning rates and camera parameter updates used in LucidDreamer [20]. Each object is allocated 1500 iterations, with the first 450 iterations serving as a warm-up. Densification and pruning begin after 100 iterations and continue until 900 iterations, with a 100-iteration interval for both operations. The azimuth angle is constrained to [-180, 180] to ensure stable optimization. The minimum timestep is 0.02, while the maximum timestep, initially 0.98, is annealed to 0.5 during warm-up and remains constant thereafter. The hyperparameters $\lambda$ and $\gamma$ in Eq. 6 are set to $8$ and $0.6$, respectively. The warm-up phase is restarted in the second stage for 30% of the remaining iterations.
A.3 Coarse Initialization Details
Each object $O_{i}$ is initialized with a point cloud $P(O_{i})$ generated using Point-E [28]. To ensure proper alignment within the scene, each point cloud undergoes a transformation based on spatial information predicted by the VLM. Let $(x_{i},y_{i},z_{i})$ denote the center coordinates, $s_{i}$ the relative size, and $\theta_{i}$ the orientation of object $O_{i}$, then the alignment transformation is given by:
$P(O_{i})_{\text{aligned}}=s_{i}\cdot R(\theta_{i})\cdot P(O_{i})+T(x_{i},y_{i},z_{i})$
where $s_{i}$, $R(\theta_{i})$, and $T(x_{i},y_{i},z_{i})$ represent scale, rotation, and translation, respectively. The full 3D scene representation is then constructed by aggregating object-aligned point clouds:
$P_{\text{scene}}=\bigcup_{i=1}^{n}P(O_{i})_{\text{aligned}}$
For independent and disentangled optimization, we track Gaussians associated with each object:
$\mathcal{G}=\bigcup_{i=1}^{n}\mathcal{G}_{o_{i}}$
where $\mathcal{G}_{o_{i}}$ contains Gaussians derived from $P(O_{i})_{\text{aligned}}$, allowing targeted refinement of object properties.
A.4 Additional Ablation Studies
Training Routine In Section 5.2, we analyze the impact of different training configurations on compositional 3D generation. Here, we provide additional results and insights. As observed in Figure 8, optimizing only relationships ($\mathcal{L}_{joint}$) produces a well-formed structure but leads to object entanglement, as boundaries between objects remain unclear. On the other hand, when relationships and objects are weighted equally ($\mathcal{L}_{joint}+\sum\mathcal{L}_{obj}$), the model struggles to properly distinguish individual elements, leading to inconsistent object representations. For instance, an astronaut riding a horse appears structurally different from a standalone astronaut, confusing the model and resulting in poor relationship modeling. Our training routine, which progressively transitions from relationship optimization to targeted object refinement, mitigates these issues. The first stage ensures that inter-object interactions are well-established, while the second stage improves individual object details while preserving these relationships.
Figure 8: Results of an ablation study on DecompDreamer under different training configurations.
Azimuth Adjustment We further analyze the effect of azimuth adjustment and object-view-dependent text prompts in multi-object scenes. Without azimuth correction, global view-dependent prompts can lead to conflicting optimizations, as different objects may have different front-facing orientations. In Figure 9, incorrect azimuth alignment results in misplaced features, highlighting the issue of optimizing objects using a fixed global view. By applying object-specific azimuth corrections, we ensure that each object is optimized relative to its correct orientation, leading to better spatial alignment and disentanglement.
Negative Prompt In Section 4.3.2, we discuss the use of negative prompts to enhance object disentanglement. Figure 10 presents experiments comparing results with and without negative prompts. For instance, in the case of the Panda, which is connected to a chair, hat, and ficus, DecompDreamer includes the Panda in the negative prompt for each connected object. As a result, there is negligible instance leakage of the Panda onto other objects. In contrast, without negative prompts, Panda fur artifacts appear on the chair. Similar effects are observed in other prompts, demonstrating the effectiveness of negative prompts in maintaining clear object boundaries.
Scene-Aware Gaussian Translation We conduct an ablation study to evaluate the impact of scene-aware Gaussian translation, as discussed in Section 4.3.1. Figure 11 compares results generated with and without Gaussian translation. Without translation, objects often fail to fully conform to their intended inter-object relationships, leading to spatial inconsistencies.
Comparison with Stable Diffusion 2.1 We further analyze the impact of using SDS-based SD2.1 versus FMD-based SD3.5 in DecompDreamer. While SDS-based SD2.1 outperforms other SOTA methods in relationship modeling, it still struggles with sharpness and convergence speed. By replacing SDS with FMD, we significantly improve object detail and reduce optimization time, requiring only one-fourth of the iterations. As shown in Figure 13, FMD produces sharper edges and more refined object details compared to SDS.
Disentanglement in Multi-Object Generation Here, we present additional results highlighting that, beyond generating compositional 3D scenes, producing high-quality, disentangled individual objects is equally critical. Figure 14 compares individual objects generated by DecompDreamer, GALA3D, and GraphDreamer on complex prompts. Both GALA3D and GraphDreamer exhibit poor disentanglement, with GraphDreamer being particularly prone to object blending due to its stronger emphasis on relationship modeling. In contrast, DecompDreamer achieves superior disentanglement through targeted relationship optimization and view-aware object refinement.
Figure 9: Results of an ablation study on DecompDreamer with (right) and without (left) the azimuth adjustment for object-view dependent text prompts.
Figure 10: Results of an ablation study on DecompDreamer with (right) and without (left) the negative prompts for the objects
Figure 11: Results of an ablation study on DecompDreamer with (right) and without (left) scene-aware gaussian translation.
Figure 12: Example usage of ChatGPT in DecompDreamer for generating compositional 3D objects.
Figure 13: More qualitative comparison of GALA3D, GraphDreamer, MVDream, LucidDreamer, Magic3D and HiFA against DecompDreamer with SD2.1 and SD3.5 as backbones.
Figure 14: More qualitative comparison of object disentanglement in GALA3D, GraphDreamer, and DecompDreamer. Each method displays the full 3D object (left) and its individual components (right).
A.5 Sample usage of ChatGPT
In Figure 12, we showcase how ChatGPT can be leveraged to generate compositional 3D objects using DecompDreamer. Given a user-defined prompt, ChatGPT generates an image and a scene graph, determining the total number of objects. It then constructs text prompts for objects, edges, and corresponding negative prompts. Next, spatial geometry is estimated, followed by 3D object initialization. Finally, the canonical 0-degree view is provided to ChatGPT to estimate the azimuth offset, ensuring accurate object orientation.
A.6 More Comparisons
In this section, we present additional qualitative comparisons of DecompDreamer against GALA3D, GraphDreamer, MVDream, LucidDreamer, Magic3D, and HiFA. Our results show that DecompDreamer outperforms existing methods by achieving more precise inter-object relationships and superior individual object modeling.
Figure 15: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20], Magic3D [21] and HiFA [44].
Figure 16: More qualitative comparison between the proposed DecompDreamer and state-of-the-art generators, including GALA3D [43], GraphDreamer [13], MVDream [36], LucidDreamer [20], Magic3D [21] and HiFA [44].